OpenStudy (anonymous):

what is the cubed root of 625 over 729?

OpenStudy (anonymous):

it is the same as the cubed root of 625 over the cubed root of 729

OpenStudy (anonymous):

\[5\sqrt{5}/9\]

myininaya (myininaya):

\[625=5^4; 729=3^6\]

myininaya (myininaya):

\[5^\frac{4}{3}=5^\frac{3}{3}5^\frac{1}{3}=5^15^\frac{1}{3}=5 \sqrt[3]{5}\] \[3^\frac{6}{3}=3^2=9\]

myininaya (myininaya):

\[\sqrt[3]{\frac{625}{729}}=\frac{\sqrt[3]{625}}{\sqrt[3]{729}}=\frac{5\sqrt[3]{5}}{9}\]

myininaya (myininaya):

the first one is what i said did you not understood what i did?

OpenStudy (anonymous):

No, would you mind explaining it? :)

myininaya (myininaya):

the best explanation i have is above is there a particular spot i lost you?

OpenStudy (anonymous):

I had trouble finding the cubed root of 729, but overall I think i understand now. Thank you.

OpenStudy (anonymous):

i would use a calculator.